how would i begin to solve a problem that is T= 2pie the square root l/g to solve for g.
Found 2 solutions by Theo, stanbon:
Answer by Theo(13342)

equation, if i understand you correctly, is:

T = 2 * pi * sqrt(1/g)

divide both sides of this equation by 2 * pi to get:

(T/(2*pi) = sqrt(1/g)

square both sides of this equation to get:

(T/(2*pi)^2 = (1/g)

multiply both sides of this equation by g and divide both sides of this equation by (T/(2*pi)^2 to get:

g = 1 / ((T/2*pi)^2)

this is equivalent to:

g = 1 / ((T^2)/(2*pi)^2) which is equivalent to:

g = (2*pi)^2 / (T^2)

note that, in general, 1/(a/b) = (1*b)/a = (b/a).

if you know the value of T, then you can find the value of g.

Assume T = 2, then, using the equation of g = (2*pi)^2 / (T^2), you will get g = 9.869604401

If you plug the value of T and the value of g in your original equation, then it should be true, assuming we converted the equation correctly.

your original equation is:

T = 2 * pi * sqrt(1/g)

replace T and g with their respective values and you get:

2 = 2 * pi * sqrt (1/9.869604401)

solve this equation to get:

2 = 2, confirming that the value for g, when T = 2, is good.


Answer by stanbon(75887)

solve a problem that is
T= 2pie the square root l/g
to solve for g
-------------------
T = 2(pi)sqrt(1/g)
----
sqrt(1/g) = T/[2(pi)]
---
1/g = [T/(2pi)]^2
----
g = [2pi/T)^2
=================